A falling body has an initial velocity of 200 cm/seconds. How far will it drop before its velocity becomes 10000 cm/seconds.


1
Expert's answer
2021-02-01T14:20:41-0500

From the Newtons third equation of motion

"v^2= u^2+ 2as"


where "v=" final velocity

"u=" initial velocity

"a=" acceleration due to gravity( 10 m/s"^2")

"s" = displacement

"100^2= 2^2+ 2 \times 9.8 \times s"

"10000 - 4= 19.6s"


"s= \frac{9996}{19.6} = 510 m"